OpenStudy (anonymous):

Build up the denominator x/x-5, ?/x2-x-20

OpenStudy (mathstudent55):

Is this what you mean? For the fractions to be equivalent, what numerator do you need in the second fraction? \(\dfrac{x}{x-5} = \dfrac{ ?}{x^2-x-20} \)

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

Start by factoring the right denominator.

OpenStudy (mathstudent55):

Do you know how to factor \(x^2 - x - 20\) ?

OpenStudy (anonymous):

Would it be x-20?

OpenStudy (anonymous):

I'm so lost! Please help!

OpenStudy (mathstudent55):

No. You need two binomials that multiply to \(x^2 - x - 20\). You need to find 2 numbers that multiply to -20 and add to -1. Can you come up with such numbers?

OpenStudy (anonymous):

-5 and 4

OpenStudy (anonymous):

Is that it? What's next?

OpenStudy (mathstudent55):

Great. Now you place the numbers like this: (x - 5)(x + 4) That's how you factor it.

OpenStudy (mathstudent55):

Now that we factored the right denominator, we have: \(\dfrac{x}{x-5} = \dfrac{ ?}{(x-5)(x+4)}\) Ok?

OpenStudy (mathstudent55):

To make the fractions equal, you need to multiply the left fraction by (x + 4)/(x + 4) \(\dfrac{x}{x-5} = \dfrac{ x(x + 4)}{(x-5)(x+4)}\)

OpenStudy (anonymous):

So the answer is x/x-5

OpenStudy (mathstudent55):

The answer is x(x + 4) which in what the right numerator needs to be for the fractions to be equivalent.